OpenStudy (anonymous):

Whats the equation for a line passing through the points (-2, -6) and (1,0)

OpenStudy (anonymous):

first find the slope y2-y1/x2-x1 then use tangent line equation of y-y1 = m(x-x1) then put it into y=mx+b m is the slope

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):